OpenStudy (anonymous):

Shelly and Terrence earned points in a game by completing various tasks. Shelly completed x tasks and scored 90 points on each one. The expression below shows Terrence's total points in the game: 90x - 20 What does the constant term of the expression represent? The total points Terrence earned The total number of tasks Terrence completed The sum of Shelly and Terrence's total points The difference between Shelly and Terrence's total points @rimshaa

OpenStudy (anonymous):

1) constant term is one without and variables/unknowns like x or y so the constant is -20. this represents that Terrence's score is 20 less than Shelly's score (90x)

OpenStudy (anonymous):

2) total points terrence earned is 90x-20

OpenStudy (anonymous):

3) total tasks completed is (if you read the question) x

OpenStudy (anonymous):

4) sum of their points would be the sum of both their equations. shelly: 90x terrence: 90x-20 sum: (90x)+(90x-20) = 90x + 90x - 20 = 180x - 20

OpenStudy (anonymous):

5) get the difference of their points by subtracting their equations. do shelly minus terrence since shelly has more points (terrence is shelly minus 20) difference: (90x) - (90x-20) = 90x - 90x + 20 (its +20 because negative and negative multiplied is always a positive) = 20

OpenStudy (anonymous):

so what would the answer be???

OpenStudy (anonymous):

for what?

OpenStudy (anonymous):

oh okay it was multiplem choice. it is the difference between their points.

OpenStudy (anonymous):

i thought you needed to find a value for each lol

OpenStudy (anonymous):

no lol

OpenStudy (anonymous):

@rimshaa so what do you think the answer is

OpenStudy (anonymous):

the answer is d. the constant term represents the difference between shelly and terrence's points